Write the slope-intercept form of an equation of the line that passes through the given point and is perpendicular to the line given. (-2,-2) y=-1/3x+9

Answer by lwsshak3(11628)

Write the slope-intercept form of an equation of the line that passes through the given point and is perpendicular to the line given. (-2,-2) y=-1/3x+9
***
form of equation for a line: y=mx+b, m=slope, b=y-intercept
y=-1/3x+9
slope,m=-1/3
..
slope,m of perpendicular line=3(negative reciprocal)
equation:y=3x+b
solve for b using given coordinates on the line(-2,-2)
-2=3*(-2)+b
b=4
equation:y=3x+4